Each of 100 firms in a perfectly competitive industry has an AVC function AVC=2Q where Q represents the output produced. If the price of the product is ₹100 find the total quantity supplied in the market when a) TFC= ₹10000 b) TFC= ₹20000


1
Expert's answer
2021-02-28T11:38:55-0500
"VC=2Q^2"

a)


"2Q^2=\frac{10000}{100}"

"Q=7"

b)


"2Q^2=\frac{20000}{100}"

"Q=10"